### 29 Mar 2026 v1

# Adapting AI to the Moment: Understanding the Dynamics of Parent–AI Collaboration Modes in Real-Time Conversations with Children

Abstract.
Parent-AI collaboration to support real-time conversations with children is challenging due to the sensitivity and open-ended nature of such interactions. Existing systems often simplify collaboration into static modes, providing limited support for adapting AI to continuously evolving conversational contexts. To address this gap, we systematically investigate the dynamics of parent-AI collaboration modes in real-time conversations with children. We conducted a co-design study with eight parents and developed COMPASS, a research probe that enables flexible combinations of parental support functions during conversations. Using COMPASS, we conducted a lab-based study with 21 parent-child pairs. We show that parent–AI collaboration unfolds through evolving modes that adapt systematically to contextual factors. We further identify three types of parental strategies—parent-oriented, child-oriented, and relationship-oriented—that shape how parents engage with AI. These findings advance the understanding of dynamic human-AI collaboration in relational, high-stakes settings and inform the design of flexible, context-adaptive parental support systems.
Human-AI Collaboration, Parent-child Interaction, Large Language Models
$c$s: Human-centered computing Collaborative interaction
Figure 1. Main screens of the COMPASS research probe: (a1) Parent pre-configuration interface for selecting collaborative functions; (a2) AI-assisted parent-child conversation interface; (b) Parent annotation interface for post-conversation annotations.
1. Introduction
Parent-AI collaboration in real-time conversational settings with children presents unique challenges. Unlike task-oriented interactions with predefined goals and structures, these settings involve sensitive and open-ended exchanges in which conversational contexts can shift rapidly within a single interaction (Driscoll et al., 2026; Morelen and Suveg, 2012; Lin et al., 2023; Overbeek et al., 2022). As a result, parents’ needs for AI support vary dynamically—for example, they may seek idea generation when fatigued (Dietz Smith et al., 2024), but expect emotional mediation in more sensitive moments (Theofanopoulou and Slovak, 2022). This variability requires parents to continuously decide when and how AI should participate, making it difficult to predefine stable collaboration strategies.
However, existing AI systems for conversational support typically rely on a limited set of static collaboration modes, such as generating prompts or acting as a conversational agent (Ho et al., 2025; Zhang et al., 2022; Liu et al., 2024; Srinivasan and Setlur, 2021). While these approaches demonstrate the potential of AI to assist in conversations, they assume that the set of collaboration modes can be predefined for the entire interaction, providing limited support for adapting AI to continuously evolving conversational contexts.
Despite the growing interest in AI-supported parenting tools, little is known about how parents’ preferred collaboration modes with AI emerge and evolve during ongoing conversations. To address this gap, we investigate how parents configure AI involvement in minimally constrained, real-time interactions. We introduce the concept of collaboration modes to describe how parents configure and engage AI scaffolding functions during interactions.
We therefore explore the following research questions:
  • RQ1: What modes of AI collaboration do parents prefer in parent-child conversations?
  • RQ2: How do contextual factors influence parents’ preferred modes of AI collaboration in parent-child conversations?
  • RQ3: What strategies do parents use to engage with AI support during parent-child conversations?
To answer these questions, we conducted a co-design study with eight parents to elicit parental needs for AI support in parent-child conversations. The study revealed six frequently mentioned functional needs (F1-F6) and two key requirements for collaborating with AI: flexible decision authority and minimal mode-switching costs.
Based on these findings, we developed COMPASS, a collaborative mediation and parental scaffolding system, as a research probe to explore parent-AI collaboration in conversation. COMPASS allows parents to flexibly combine parental support functions during conversations, enabling us to examine how their strategies adapt to shifting collaboration needs.
Using COMPASS as a research probe, we conducted a lab-based study with 21 parent-child pairs in semi-structured conversational tasks. We found that parents developed preferred collaboration modes with AI, which moved towards a more partner-like collaboration mode and converged into recurring patterns (RQ1). Parents adjusted these modes systematically based on emotional intensity, emotion type, conversation stage, parental goals, and conversation scale and structure (RQ2). Additionally, parents adopted distinct strategies to engage with AI support across three dimensions: parent-oriented, child-oriented, and relationship-oriented (RQ3).
Our study contributes to HCI in four ways. First, we provide empirical insights into the dynamic nature of parent-AI collaboration, showing how collaboration modes evolve over time, vary with contextual factors, and are enacted through parental strategies. Second, we extend Parental Goal Theory by grounding its three established categories in our empirical findings of AI-assisted parenting. Third, we advance the understanding of human-AI collaboration in relational, high-stakes domains. Finally, we offer design implications for parental support systems that enable flexible, context-adaptive human-AI collaboration.
2. Related Work
2.1. Parent–Child Conversations in HCI
Parent–child communication is critical to early childhood development, supporting children’s cognitive, social, and emotional growth through everyday interactions (Vygotsky, 2012; Hart et al., 1997; Grotevant, 2001; McLean et al., 2007; Carson et al., 1999; Fitzpatrick and Koerner, 2005). As children’s first social environment, families position parents as primary agents of socialization. In these interactions, parents continuously guide, respond to, and adapt to children’s needs, shaping the quality and outcomes of communication.
Prior HCI research has explored supporting parent–child conversations from multiple perspectives (Zapf et al., 2023), including (1) content provisioning (e.g., prompts or games) (Chan et al., 2017); (2) process guidance that analyzes conversation dynamics and provides feedback (e.g., tone or emotional cues) (Choi et al., 2025; Dietz Smith et al., 2024); and (3) third-party mediators (e.g., social robots) that restructure interaction (Ho et al., 2024, 2025; Seo et al., 2025). While these approaches provide valuable support, they largely assume stable interaction patterns, limiting adaptation to dynamically evolving conversational needs.
In practice, parent-child conversations are highly dynamic and context-sensitive. Parents often balance multiple, shifting goals—such as guidance, protection, reciprocity, and control—while responding to children’s changing cognitive, emotional, and attentional states (Krevans and Gibbs, 1996; Knafo and Schwartz, 2003; Grusec and Goodnow, 1994; Morelen and Suveg, 2012; Denham, 1998; Flavell, 1994; Posner and Rothbart, 2007). This requires parents to continuously adjust how they engage in conversations.
Prior work has explored parent–AI collaboration in conversational settings (Ho et al., 2023; Zhang et al., 2022). However, how such collaboration evolves in response to changing goals and contexts has not been systematically examined. Our work addresses this gap by investigating the dynamics of parent–AI collaboration in real-time conversations and showing how these dynamics can inform the design of adaptive AI support for parent–child communication.
2.2. AI Systems for Parent-Child Interaction
Prior AI systems for parent-child education typically predefine parent-AI roles before interaction. Existing work identifies three collaboration modes. In the parent-led, AI-assisted mode, AI supports parents’ emotional guidance, story co-creation, and content selection through analysis, generation, and feedback (Shen et al., 2025; Xu et al., 2025; Kim et al., 2025; Zhao et al., 2025). In the AI-led, parent-assisted mode, AI acts as the primary interactive agent while parents supervise or intervene when needed (Lau et al., 2025; Zhang et al., 2025a). A third mode emphasizes joint collaboration, positioning AI as a conversational partner in multi-turn interactions (Viswanathan et al., 2025; Shang et al., 2025).
While these systems explore different role allocations, they primarily focus on static collaboration structures (Li et al., 2025; Wang et al., 2026), with less attention to how parents can adapt AI involvement during ongoing interactions. In practice, parent-child conversations are highly dynamic, with emotions, goals, and engagement shifting in real time (Morelen and Suveg, 2012; Bradley et al., 2015). Supporting parents in responding to these changes requires more in situ control over AI behavior.
Recent work has begun to explore more flexible forms of parental participation. For example, parents express a desire to adjust AI content generation, interaction depth, and collaboration structures according to contextual factors such as children’s learning stages, time availability, and energy levels (Ho et al., 2025; Zhang et al., 2022; Ho et al., 2023, 2024). Other studies introduce configurable parental control tools that support differentiated permissions, interaction modes, and intervention intensity across developmental stages (Dumaru and Al-Ameen, 2025; Yang et al., 2023).
However, these approaches largely assume relatively static collaboration modes throughout the interaction, offering limited support for adapting AI behavior to rapidly changing conversations. In addition, prior work provides limited theoretical grounding for understanding how and why parents adjust collaboration with AI in real time. Parental Goal Theory (Hastings and Grusec, 1998) suggests that parents simultaneously pursue parent-, child-, and relationship-oriented goals that shift during interaction. Building on this perspective, we extend the theory to explain how parents regulate collaboration with AI under changing contextual factors.
2.3. Collaboration Dynamics in Human-AI Interaction
HCI research has long examined how humans and AI share roles in collaboration. Prior work often frames AI through static roles. As a tool or assistant, AI improves efficiency while humans retain primary control (Wu et al., 2022; Wang et al., 2019a; Lee et al., 2022). As a coach or mentor, AI provides guidance without taking over decisions (Arakawa and Yakura, 2024; Fadhil et al., 2019). As a teammate or partner, AI collaborates with users toward shared goals through complementary capabilities (Zhang et al., 2021; Brandt et al., 2009; Salikutluk et al., 2024). However, these roles are typically fixed at design time, limiting collaboration to stage-based arrangements and overlooking interaction dynamics.
Recent work instead frames collaboration as “agency negotiation”, where initiative and control shift with task stages and contextual factors (Masson et al., 2024). Adjustable autonomy has been shown to improve perceived usefulness, competence, and satisfaction (Houde et al., 2025; Zhang et al., 2023), and to enhance team performance when AI adapts during collaboration (Salikutluk et al., 2024). While prior work has begun to highlight interaction-level dynamics, most systems examine these dynamics at coarse-grained task units (e.g., storytelling or search sessions) (Zhang et al., 2022; Sharma et al., 2024), with limited work systematically capturing fine-grained interaction changes or validating them through system design.
Recent systems explore finer-grained collaboration mechanisms, including proactive intervention timing (Oh et al., 2024; Kuang et al., 2024; Zhu et al., 2025), proactive task collaboration (Pu et al., 2025; Mei et al., 2025), shared representations for alignment (Jeong et al., 2025; Zhang et al., 2025b), and role fluidity at the micro-task level (Zhou et al., 2024; Liu et al., 2025; Wu et al., 2025).
However, these studies primarily address general task collaboration and rarely examine parent-child communication, a setting characterized by multiple concurrent goals (Krevans and Gibbs, 1996; Knafo and Schwartz, 2003), strong emotional fluctuations (Denham, 1998), and asymmetric roles (Elbers*, 2004). Moreover, they lack a theoretical account explaining how users select collaboration modes under such conditions.
Building on this literature, we examine collaboration dynamics in emotionally sensitive parent–child conversations, tracing how parents adjust control and support in real time and extending prior frameworks from macro-level autonomy to micro-level interaction.
3. Formative Study
To understand parents’ needs and preferred modes of collaborating with AI in parent–child conversations, we conducted a formative co-design study to inform the design of our research probe.
3.1. Setup
3.1.1. Protocol
The study consisted of a pre-task and a co-design session. Before the session, parents engaged in a 10-minute conversation with their child about daily challenges (e.g., peer conflict, teacher criticism), audio-recorded the interaction, and shared it with the research team.
During the co-design session, parents reviewed their recordings with two HCI researchers and identified moments where AI support could be helpful using an online whiteboard (Figure 2). Conversations were structured into four phases (before, during, after, and throughout), and parents annotated desired AI support for each phase. They then organized, prioritized, and combined these functions, and discussed feature-switching preferences and usage requirements. Participants were encouraged to think aloud, and facilitators used follow-up questions to elicit detailed reflections.
Figure 2. Illustration of the co-design session.
3.1.2. Participants
We recruited eight parents (P1-P8) of children aged 6-8 via social media. This age group was selected because children in the early primary school years still rely on parental guidance (Sroufe et al., 2005; Gottman et al., 2013) while developing conversational abilities that allow them to engage in meaningful conversations (Berman and Slobin, 2013; McCarthy, 1943).
Participants included eight parents (5 fathers, 3 mothers; $M=38.62$, $SD=4.18$) and their children (2 boys, 6 girls; $M=6.50$, $SD=0.71$). Sessions were conducted via video conferencing ($\sim$1 hour) and facilitated by two HCI researchers. All participants provided informed consent and received \$10 or equivalent compensation.
3.2. Findings
3.2.1. Scaffolding Functions Aligned with Parental Goals
From the co-design sessions, we identified six frequently requested AI scaffolding functions (F1-F6), which we group into two categories based on Parental Goal Theory (Dix, 1991).
Parent-focused goals include Scenario Provision (F1), Utterance Suggestion (F2), and Reflective Feedback (F3), supporting parents in framing topics, expressing themselves, and reflecting on their strategies. Parents preferred contextualized storytelling over fixed scripts (F1; P6:“making it easier for children to understand and accept”), concise and directly usable utterances at critical moments (F2; P8:“could read aloud directly”), and non-judgmental, heuristic feedback (F3; P2:“I need it to inspire me on how to think”).
Relationship-focused goals include Conversation Summarization (F4), Agent-as-Speaker (F5), and Multimodal Enrichment (F6), supporting mutual understanding and engagement. Parents expected real-time summaries to track emotions and perspectives (F4; P1:“so that everyone can understand each other”), occasional AI participation as a lightweight third party when conversations stalled (F5; P4:“third-party friend”), and multimodal content (e.g., images) to help children understand the current conversation and sustain engagement (F6; P5:“he will be more willing to continue the conversation”).
Child-focused goals were excluded, as our study focused on parent-centered collaboration needs in conversations. Details of the six functions (F1–F6), including their purposes and use contexts, are provided in Appendix A.
3.2.2. Flexible Function Combination with Retained Decision Authority
In the co-design sessions, parents emphasized the need to dynamically adapt scaffolding functions to rapidly changing conversational contexts.
Preferences varied. Some parents preferred full manual control, advocating “free combination” to maintain flexibility and natural interaction: “The freer, the better… the more fixed its feedback format is, the less natural it feels.” (P4). Some parents favored preconfigured function combinations to reduce effort: “I think fixed system functions would be better, so I don’t have to switch frequently.” (P7)
Most parents preferred a hybrid approach, where AI suggests an initial set of combinations while leaving final control to parents. P6 noted, “You can offer suggestions for them to adopt,” and P8 suggested “it could prompt whether I need certain functions and I can just select ‘yes’.
Across these views, parents consistently stressed retaining decision authority in choosing functions (P4: “Different parents may have different needs”), as needs vary across situations and conversations are inherently unpredictable (P8).
3.2.3. Minimizing Interactional Costs of Collaboration Mode Switching
Parents noted that frequent manual switching could disrupt conversational flow, suggesting a preference for minimally intrusive interaction (P7: “when talking to children…keeping a fixed mode might be better”).
To reduce switching costs, parents proposed more seamless interaction modalities. Voice control was seen as effective when typing is inconvenient (P5: “Voice commands… more convenient and frees my hands”), while many favored lightweight, intuitive interfaces such as one-click toggles (P2: “Clicking is best—it matches our usual habits.”). Some also suggested drag-and-drop interactions for flexible configuration (P6).
Overall, parents preferred switching mechanisms that are low-interference, predictable, and easily reversible, enabling real-time adaptation without disrupting conversational continuity.
4. COMPASS Design and Implementation
Building on insights from the formative study, we designed and implemented COMPASS as a research probe to support flexible parental scaffolding in parent–child conversations. Rather than functioning as a full intervention system, COMPASS is intended to investigate how parents’ collaboration needs and behaviors with AI evolve during such conversations.
To preserve natural decision-making, we minimized system-induced bias that could influence parents’ collaboration choices.
To support analysis of parent-AI collaboration dynamics, we incorporated an annotation interface to capture in-situ reflections and strategies.
Figure 3. Overview of the COMPASS research probe workflow.
4.1. User Interface Design
The COMPASS interface consists of two components: Parent–Child Conversation and Parent Annotation (Figure 1).
4.1.1. Parent-Child Conversation Interface
This interface helps parents conduct topic-based conversations with their children. Based on the findings from the formative study, we derived a set of design requirements (Table 1).
Table 1. Mapping from formative study findings to COMPASS design features.
Formative Study Finding COMPASS Design Feature
Scaffolding functions aligned
with parental goals
Six modular functions in the
Parent-Child Conversation Interface
Flexible function combination
with retained decision authority
Initial function set based on stated purpose;
parents can enable, disable, or reorder at any time.
Minimizing interactional costs of
collaboration mode switching
Auto-update the active modules every 30 s;
enable/disable functions with a single click
To address the identified functional needs (F1–F6), COMPASS includes six modules: Scenario Provision, Utterance Suggestion, Reflective Feedback, Conversation Summarization, Agent-as-Speaker, and Multimodal Enrichment. Each module provides contextual support during the conversation.
Before the conversation, parents configure an initial set of functions (a1 in Figure 1). They can describe their communication needs in natural language, based on which COMPASS suggests a set of functions via LLM-based reasoning. Parents can also manually adjust the collaboration mode.
During the parent–child conversation (a2 in Figure 1), COMPASS provides support through the enabled modules. Parents can enable or disable modules at any time, enabling flexible adaptation to evolving interactions. To balance responsiveness and minimal disruption, continuous-support modules (F1-F4) are automatically updated every 30 seconds when enabled, while the remaining modules (F5-F6) are triggered on demand. Functional modules are presented in a compact list, enabling quick switching with minimal cognitive and operational effort.
4.1.2. Parent Annotation Interface
To examine the rationale behind parents’ collaboration choices, COMPASS includes a post-conversation annotation interface (b in Figure 1). The conversation is segmented into slices based on moments when parents change the activation of one or more modules; consecutive changes are merged into a single slice if no conversation occurs between them.
For each conversation slice, COMPASS presents the transcript, interface snapshot, and audio replay to support recall. Parents are then asked to respond to structured questions about contextual factors during that slice, including their own state, the child’s state, and interaction-related factors (see all rows labeled “Annotator = Parent” in Table 2). An example of a single annotation is shown in Figure 4.
Figure 4. Example parent annotation for a conversation slice.
4.2. COMPASS Workflow
We describe the workflow of COMPASS as a research probe supporting parent–child conversations (Figure 3).
During the parent-child conversation, the Audio Processor performs automatic speech recognition (ASR) and speaker diarization on the live audio stream. The resulting transcripts, annotated with speaker labels, are passed to the Conversation Monitor, which maintains a sliding window of recent utterances.
Meanwhile, parents can enable or disable functions (F1-F6) through the Parent–Child Conversation interface, which updates the Function State Manager in real time. Parents may select a predefined topic from the Scenario Manager to initiate a conversation. The scenario set includes eight situations identified in the formative study as challenging for children aged 6-8 (e.g., peer conflict).
The current context, including recent utterances, function states, and the selected scenario, is passed to the Function Library to generate context-aware support for the parent. The Function Library contains six modules (Section 3.2.1) that operate independently. These modules leverage LLMs, text-to-speech, and text-to-image models to provide multi-channel support.
Implementation details are provided in the Appendix B.
5. User Study
We conducted a lab-based study with 21 parent-child pairs using COMPASS as a research probe to examine (1) parents’ preferred AI scaffolding modes, (2) how conversational contextual factors shape these preferences, and (3) strategies parents use to engage with AI during parent–child conversations. The study was IRB-approved, and informed consent/assent was obtained from all participants.
5.1. Participants
We recruited 21 parent-child pairs via social media. Each pair received \$20 or equivalent compensation. Parents (P1-21) included 4 fathers and 17 mothers ($M=39.05$, $SD=4.02$). Children (C1-21) included 11 boys and 10 girls aged 6-8 ($M=7.00$, $SD=0.76$); the rationale for this age group is provided in Section 3.1.2. Detailed demographics are provided in Appendix C.
All sessions were recorded (video, audio, screen) with anonymization. Throughout the study, system outputs were monitored remotely to prevent harmful content.
5.2. Procedure
Each parent-child pair participated in a continuous two-hour lab session using a touchscreen laptop.
Introduction (25 min). Parents were introduced to COMPASS and its six modular functions, followed by a 10-minute free conversation to familiarize themselves with the system. They were encouraged to explore different feature combinations and collaboration strategies.
Parent–child conversation (20 min). Parents and children discussed one of two randomly assigned challenging topics (e.g., peer conflict, teacher criticism), selected based on formative interviews with five parents. Parents interacted with the system while guiding their children.
Co-annotation (45 min). Using the Parent Annotation Interface, parents and an HCI researcher jointly reviewed conversation segments. Parents recalled interactions via replay, annotated contextual factors (Table 2), and explained their decisions through researcher probing. Children stayed in a supervised adjacent area during this and the next session.
Final interview (15 min). We conducted semi-structured interviews to elicit parents’ overall strategies for engaging with AI.
5.3. Data Analysis
For quantitative analysis, we pre-processed COMPASS logs and questionnaire data. Numeric features (e.g., turn count, utterance length) were z-standardized. For each conversation slice, collaboration modes were encoded as six-digit binary vectors indicating the activation of each function. Questionnaire responses (e.g., stage, emotion type) were one-hot encoded. We then analyzed function usage over time and its correlations with contextual factors.
For qualitative analysis, we transcribed co-annotation and interview sessions and conducted thematic analysis (Braun and Clarke, 2006; Terry et al., 2017; Braun and Clarke, 2019). Four HCI researchers performed iterative coding: initial open coding, collaborative codebook development with one participant (P1), and independent coding with consensus discussions. Codes were then clustered to identify recurring parental strategies and collaboration patterns. All analyses were conducted in Mandarin to preserve original meanings.
6. Findings
6.1. Parents’ Preferred Modes of AI Collaboration (RQ1)
Figure 5. Overall function enable rates (95% CI).
Figure 6. Function enable rates across the first and second halves of participants’ sessions.
6.1.1. Moving Towards a More Partner-like collaboration mode
Figure 6 shows that Utterance Suggestion (F2) and Conversation Summarization (F4) were most frequently used, while Multimodal Enrichment (F6) was least used; Scenario Provision (F1), Reflective Feedback (F3), and Agent-as-Speaker (F5) showed moderate usage.
Temporal analysis (Figure 6) revealed a significant increase in Multimodal Enrichment (F6, $132.5\%$, $p<0.01$), with smaller gains in Agent-as-Speaker (F5, $28.9\%$) and Reflective Feedback (F3, $17.0\%$), and declines in Scenario Provision (F1, $19.6\%$) and Conversation Summarization (F4, $10.4\%$).
Overall, parents’ preferences shifted toward a more partner-like collaboration mode, characterized by reduced reliance on structural scaffolding (F1, F4) and increased engagement in co-performance with AI (F5, F6).
Figure 7. Distribution of the eight most frequent collaboration modes observed in our study.
Figure 8. Heatmap of the correlations between each pair of functions.
6.1.2. Adding Reflective or Expressive Support on Top of Basic Assistance
Parents explored 41 out of 64 possible collaboration modes, indicating diverse and flexible needs beyond fixed configurations. Pairwise correlations among functions were weak ($\phi\in[-0.2,0.2]$; Figure 8), suggesting that functions were combined largely independently.
Despite this diversity, usage was highly skewed: the top eight modes accounted for 51.20% of all cases (Figure 8). These modes reveal consistent structural patterns:
Basic assistance as a constant backbone. Utterance Suggestion (F2) and Conversation Summarization (F4) were always enabled (8/8), with Reflective Feedback (F3) frequently co-occurring (6/8).
Reflection as compensation. When Scenario Provision (F1) was absent, Reflective Feedback (F3) was consistently used (3/3).
Expressive support as co-performing signal. Multimodal Enrichment (F6) only appeared with Agent-as-Speaker (F5) (3/3), indicating that richer expression was adopted when AI took a more active role.
Overall, parents anchored on basic assistance (F2, F4) and selectively layered reflective (F3) or expressive support (F5, F6), revealing a structured yet flexible strategy for configuring collaboration.
6.2. Contextual Factors Influence Collaboration Modes (RQ2)
We examined how contextual factors shape parents’ collaboration modes, drawing on Parental Goal Theory (Dix et al., 1986; Hastings and Grusec, 1998). Specifically, we focus on contextual factors that are observable in lab settings and have been widely discussed in prior work. Factors were grouped into parent-, child-, and relationship-oriented categories (Table 2). Correlation analysis (Figure 9) shows that function usage systematically varies with contextual factors.
Table 2. Contextual factors explored in this study. The citations in the “Data Type” column refer to prior works that discuss these contextual factors.
Category Contextual Factor Annotator Data Type
Parent- oriented Cognitive Load Parent Numeric (Hart and Staveland, 1988)
Emotion Type Parent Single-choice (Ekman, 1992)
Emotion Intensity Parent Numeric (Roseman, 1991)
Energy Level Parent Numeric (Hart and Staveland, 1988)
Speaker Turns AI Numeric
Speaker Length AI Numeric
Child- oriented Emotion Type Parent Single-choice(Ekman, 1992)
Emotion Intensity Parent Numeric (Roseman, 1991)
Speaker Turns AI Numeric
Speaker Length AI Numeric
Relationship- oriented Conversation Stage Parent Single-choice(Mercer and Littleton, 2007; Graesser et al., 2005)
Conversation Feature Parent Single-choice (Mercer and Littleton, 2007)
Conversation Goal Parent Single-choice (Wood et al., 1976)
Total Turns AI Numeric
Conversation Length AI Numeric
Figure 9. Heatmap of correlations between functions and contextual factors, using point-biserial (binary–continuous) and Phi (binary–binary) coefficients. Statistical significance is indicated by asterisks (***: $p<0.001$, **: $p<0.01$, *: $p<0.05$).
6.2.1. Parents’ States Lead to Diverse Collaboration Modes
Parents’ internal states strongly shape AI roles.
Cognitive load and energy determine support level. Under high cognitive load, parents prefer direct scaffolding, increasing use of Utterance Suggestion (F2, $p<0.001$) and Conversation Summarization (F4, $p<0.01$). When energy is low, they avoid complex support such as Multimodal Enrichment (F6, $p<0.05$).
Emotional intensity shifts AI from guide to mediator. Higher parent emotion intensity reduces acceptance of directive support (F1, F2) but increases reliance on Conversation Summarization (F4) and Agent-as-Speaker (F5) ($p<0.05$), indicating a shift toward neutral mediation or delegation.
Perceived helpfulness enables deeper collaboration. Higher trust is associated with increased use of reflective and summarizing support (F3, F4; $p<0.05$), reflecting a transition from using AI as a “tool” to collaborating with it as a “partner”.
6.2.2. Children’s Emotional States Lead to Polarized Collaboration Modes
Children’s emotional states induce a polarized pattern.
Calm and positive states encourage empowerment. When children are calm or positive, parents favor Scenario Provision (F1, $p<0.001$), Multimodal Enrichment (F6, $p<0.01$), and reflection (F3, $p<0.05$). This suggests that in low-stress, open communication environments, parents have more cognitive resources to interact with AI support, such as reference cases or multimedia elements.
High-intensity negative states reduce directive guidance. When children are sad, angry, or highly aroused, parents decrease use of F1 and F2 ($p<0.01$), avoiding prescriptive suggestions.
High emotion intensity shifts AI to front-line support. With increasing emotion intensity, parents rely more on Conversation Summarization (F4) and Agent-as-Speaker (F5), using AI as a neutral mediator or substitute speaker to share their burden.
6.2.3. Relationship-oriented Factors Reshape AI Roles
Parents’ collaboration choices are shaped by relationship-oriented factors, including conversation stage, goal, and interaction intensity.
Conversation stage shapes AI support. Parents rely on AI for sensemaking in the early Problem Identification stage (more F4, less F5; $p<0.05$) and for idea generation in the Exploration stage (more F2; $p<0.001$), but reduce reliance on content suggestions in the Emotion Expression and Problem Solving stages (less F2; $p<0.01$).
Conversation goals shape whether AI speaks directly. Under the Regulate goal, parents reduce F1/F2 ($p<0.001$) and increase Agent-as-Speaker (F5; $p<0.05$), delegating persuasion to AI. Under the Highlight goal, they increase F1/F2 (p<0.05p<0.05/$p<0.001$), using AI as a behind-the-scenes supporter to strengthen their own voice.
Interaction intensity shapes AI roles. As conversation turns increase, parents use more Conversation Summarization (F4; $p<0.05$) and less Agent-as-Speaker (F5; $p<0.05$), positioning AI as an organizer of the conversation rather than a third speaker.
6.3. Three Types of Strategies Parents Employ to Engage with AI (RQ3)
Based on interview data, we identify three strategy types guiding parent-AI engagement: parent-oriented, child-oriented, and relationship-oriented. These extend Parental Goal Theory (Dix et al., 1986; Hastings and Grusec, 1998) to AI-supported parent-child interaction.
6.3.1. Parent-Oriented Strategies
Parents adopted self-focused strategies to (1) compensate for their limitations and (2) regulate their own cognitive and emotional states.
(1) Compensating for limitations and maintaining conversational control. 11 of 21 parents used AI for real-time feedback and information retrieval to compensate for communication limitations, with eight reporting that AI helped identify deficiencies and support self-correction. As P11 noted, “it can point out some parts of my speech that are not quite right… I can adjust my direction accordingly.” Parents also relied on AI to address knowledge gaps (e.g., explaining “new moon” to children) and to provide psychologically appropriate responses. In emotionally sensitive situations, AI’s phrasing suggestions were used to avoid potentially harmful expressions.
9 parents further leveraged AI to improve interaction quality. Seven described using emotional insights to “shift from evaluative to empathetic communication” (P18), while six relied on AI’s suggestions to maintain conversational flow when unsure how to initiate a topic (P3).
Under high stress or uncertainty, some parents adopted an “all-round reliance” strategy by enabling multiple functions simultaneously to maintain conversational control. As P4 noted, “I was worried that after my child expressed something, I would not know how to respond. So I kept Functions 2, 3, and 4 enabled to make sure that after my child spoke, I could immediately pick up her words.
(2) Managing parents’ emotional and cognitive states. Parents reported using AI to regulate their internal states during conversations. Three described using reflective prompts to manage agitation and prevent negative emotions (P14: “parents’ emotions can affect the conversation, and AI can help me with that”). When distracted by work or other demands, two parents used summarization and prompting to refocus and maintain continuity, noting that the system could “help pull me back to the core of the conversation” (P7). Additionally, parents used AI to broaden thinking when constrained by habitual patterns, seeking alternative perspectives and inspiration (P2: “When I feel like my words are confined to my habitual ways of thinking, I try (using AI) to get some help”).
6.3.2. Child-Oriented Strategies
Parents adapted AI use by continuously observing and supporting their children’s behaviors, focusing on (1) monitoring children’s status and (2) addressing children’s expressive difficulties.
(1) Observing and evaluating children’s short- and long-term status. Parents dynamically adjusted system features based on children’s attention. For instance, some enabled or avoided multimodal features depending on whether they supported or distracted engagement, noting that “the child might start talking to the AI, which would distract his attention” but could also “raise his interest” (P4).
Parents also relied on system feedback and scenario-based prompts to better perceive children’s emotions. They report that system feedback can help them perceive their child’s feelings more sensitively. Some parents further highlight that Scenario Provision) (F1) allows them to observe the child’s emotions and understand their “first reaction” and “true thoughts” (P4).
Beyond moment-to-moment interaction, parents considered long-term factors such as age, interests, and personality. They adjusted feature use accordingly—for example, reducing “childish” cases for older children (P9) or leveraging familiar scenarios to help children “immerse more easily” (P3) and “better control their emotions” (P19). These practices reflect a child-adaptive strategy of tailoring AI support to developmental and individual needs.
(2) Coping with children’s expression barriers. When children showed resistance or struggled to express themselves, parents used AI as a “backup guide” to sustain interaction. They enabled Utterance Suggestion to restart stalled conversations (P3) or used multimodal cues to redirect attention (P1). Parents also leveraged AI to regulate conversational pacing. Several parents highlighted that Reflective Feedback (F3) is particularly effective in slowing down the interaction. Pauses for self-reflection “give the child more space to organize their language” (P7).
To stimulate expression, parents employed AI-generated cases, prompts, or scenarios to elicit responses. For example, reading AI-suggested cases helped children “align their thinking” (P4), while prompts encouraged children “be willing to carry the conversation more deeply” (P6). In some cases, multimodal content was used at the end of conversations to foster emotional resonance and extend thinking (P4).
6.3.3. Relationship-Oriented Strategies
Beyond parent- or child-focused strategies, parents positioned COMPASS as a relational mediator that both directly intervenes in conversation and indirectly shapes interaction.
(1) Direct AI involvement as a mediator. Parents used AI as a neutral third party to prompt children’s expression and reinforce their own perspectives. For example, AI could “guide the child to express inner thoughts, just like a psychological expert” (P19). AI could also repeat parental viewpoints (P17:“it made my opinion more convincing to the child ”).
Some also relied on AI to sustain interaction, keeping it active to facilitate turn-taking and help children elaborate incomplete responses (P4:“ Adding Function 5 helped inspire my child to say more and finish their sentences ”).
AI further functioned as an emotional buffer in tense situations, helping both sides regulate emotions and continue communication. As one parent put it: “The biggest function (of COMPASS) is to control emotions.” (P19).
(2) Indirect influence on relationship dynamics. Parents described COMPASS as shaping interaction climate through subtle, indirect influence. Its presence encouraged more serious and reflective engagement (P17: “it can help both my child and me take (this topic) more seriously”).
AI-generated cues also fostered moments of closeness and responsiveness. Parents reported that AI-generated prompts helped foster intimacy in ways they might not have achieved on their own. For example, after COMPASS generated an illustration for a storybook, one parent recalled: “My child sat on my lap, just like in the AI storybook, and we had a close physical contact. ” (P1).
Finally, parents used AI to check whether they and their children were “on the same page.” Real-time summaries from F4 reassured them about alignment: “I want to check (whether) our communication is (happening)… (whether) we are engaging in effective communication” (P20).
7. Discussion
7.1. Evolving Dynamics of Parent-AI Collaboration with COMPASS
Our findings reveal how parent–AI collaboration unfolds over time. We highlight three representative themes that extend prior work and offer new design insights.
7.1.1. From Backstage Support to Partner-style Co-performance
Parents initially relied on backstage support to maintain conversational flow, then gradually granted COMPASS more visible roles (F5, F6), moving toward partner-style collaboration. Extending prior work on flexible role division (Zhang et al., 2022; Gvirsman and Gordon, 2024; Ho et al., 2025), we identify a risk-perception-driven shift: parents incrementally delegate control to assess AI controllability before incorporating it into the “safe zone” of parent–child interaction. This suggests that systems should support progressive trust-building rather than assuming immediate readiness for partnership. It extends HCI trust research into parenting contexts (Niu et al., 2023; Chen et al., 2024b; Jo et al., 2025).
7.1.2. From Experimentation to Stable Collaboration Modes
In the user study, only a subset of collaboration modes was consistently used, while others appeared as short-lived transitional states. Beyond prior work treating collaboration styles as stable preferences (Ho et al., 2024, 2023), our findings show that stable and transitional modes coexist. Transitional modes reflect parents’ experimentation and familiarization with the AI system, echoing exploratory processes in other HCI domains (e.g., exploratory search) (Klouche et al., 2015; Sarma et al., 2024). This suggests that systems should scaffold low-risk experimentation to help families converge on stable strategies suited to their conversational settings.
7.1.3. Context-Aware Restraint in Sensitive Parenting Moments
Parents adapted AI use to context: in calm, cognitively oriented moments, they preferred structured support, whereas in emotionally intense situations, they avoided directive scaffolds and relied on alignment and mediation. While prior HCI work highlights context-aware scaffolding (Ogata and Yano, 2004; Chen et al., 2023, 2025b), we further identify the need for context-aware restraint. AI should not only know when to step in but also hold back in “no-go” moments (e.g., avoiding rigid responses when children are sad). This aligns with “calm technology” (Weiser and Brown, 1996, 1997) and reframes deliberate AI silence as a form of positive intelligence that protects sensitive parent–child interactions.
7.2. From Parental Goal Theory to AI-Engaged Parental Goal Theory
Parenting goals are typically classified as parent-oriented, child-oriented, and relationship-oriented goals (Dix et al., 1986; Hastings and Grusec, 1998). Grounded in our findings, we revisit these categories and show how they are reshaped in AI-engaged parenting contexts.
7.2.1. Orchestrating AI Support for Parent-Oriented Goals
Prior work shows that parents aim to maintain competence, bridge knowledge gaps, and regulate their own states when interacting with children (Coleman and Karraker, 1998; Deater-Deckard, 1998; Heerman et al., 2017). Our findings support this view, but further show that these goals are increasingly achieved through AI collaboration rather than solely individual effort. Parents treated AI as a just-in-time collaborator that compensated for gaps in knowledge and experience, supported conversational flow, and reduced uncertainty in stressful situations. When facing high cognitive load or emotional instability, they dynamically adjusted collaboration modes to obtain appropriate support, often drawing on familiar technology-use habits. This suggests a shift from individual regulation to parent-AI collaborative regulation in parent-oriented goals.
7.2.2. Extending AI-in-situ Feedback to Support Child-Oriented Goals
Child-oriented goals are typically framed around long-term cognitive and socio-emotional development (Chang and Lee, 2017; Dix, 2014; Ho et al., 2022; Sun et al., 2024), and existing AI systems largely follow this orientation by supporting children’s knowledge learning, emotion recognition, and self-expression (Sun et al., 2024; Chen et al., 2024a, 2025a).
Our findings extend this view by showing that child-oriented goals also involve attending to children’s moment-to-moment states (e.g., emotional type, emotional intensity, and expressive state), which is underexplored in prior theory and system design. While parents are known to regulate children’s emotions during parent-child interactions (Eisenberg and Spinrad, 2004; Eisenberg and Morris, 2002), we find that AI reshapes this process: parents rely on AI to identify and summarize children’s current states, such as emotional type and emotional intensity, reducing the cognitive burden of continuously monitoring children’s states during parent–child conversations.
By activating state-reflective AI features, parents gained a clearer understanding of children’s moment-to-moment states, which directly informed their selection of AI collaboration modes. This suggests that AI expands child-oriented goals from a sole focus on long-term development to a combination of developmental goals and moment-to-moment regulation. Future work should examine how AI can support and balance these two aspects without displacing parental judgment.
7.2.3. AI-Mediated Parent-Child Conversation to Support Relationship-Oriented Goals
Prior work frames relationship-oriented goals as maintaining and enhancing the quality of parent–child relationships (Laursen and Collins, 2009, 2003). Our findings extend this view by showing that AI actively reshapes relational dynamics, as it supports interaction continuity, encourages children’s disclosure, and buffers emotional tension. This is particularly important under high emotional load, where parents often struggle with real-time relationship regulation, leading to escalation and conflict (Barros et al., 2015; Morris et al., 2007).
With AI introduced as a new participant in parent–child conversations, relational maintenance is no longer solely the parent’s responsibility. Parents positioned AI as both a conversational mediator and a relational scaffold: as a neutral third party that encouraged children’s disclosure and helped de-escalate tension, and as a support for alignment and intimacy within the interaction.
This transforms relationship-oriented goals from a dyadic parent–child structure to a triadic parent–child–AI collaboration, extending the theoretical boundary of relational goals. Moreover, parents’ trust and attitudes toward AI shaped how much relational authority they delegated, requiring a balance between AI participation, parental authority, and children’s autonomy. Together, these findings suggest that AI functions as a collaborative social actor in family interaction, fundamentally reshaping the structure of relationship-oriented goals.
7.3. Towards More Fine-grained and Dynamic Human-AI Collaboration Framework
Prior research on human-AI collaboration largely focuses on task-driven contexts, where success is defined by performance metrics such as accuracy and efficiency (Wang and Yin, 2022; Chattopadhyay et al., 2017; Taesiri et al., 2022). Such work often assumes that collaboration goals are clearly defined and relatively stable, and that the division of labor between humans and AI can be specified in advance of interaction. Recent studies introduce human-centered factors (e.g., trust, transparency) to collaboration frameworks (Duan et al., 2025; Vössing et al., 2022), yet often treat these goals in isolation, overlooking how multiple goals interact in practice. Our findings extend this view, showing that parent–AI collaboration in our conversation setting involves navigating multiple, sometimes conflicting goals.
In parent-child conversations, parents simultaneously balance three goals: parent-oriented, child-oriented, and relationship-oriented. The priority of these goals shifts across situations, and parents’ ongoing trade-offs across these goals lead to more fine-grained and dynamic parent-AI collaboration. This goes beyond prior work, which assumes a stable collaboration mode throughout interaction (Ho et al., 2025; Zhang et al., 2022).
In our study, parents switched collaboration modes 7.90 times per conversation on average. With only six functions, they produced 41 of 64 possible combinations, indicating that collaboration modes are not predefined but constructed in situ. Through frequent switching and recombination of AI functions, parents continuously adjust AI’s role, form, and level of involvement in response to conversational progress, the child’s state, and their own judgments. These results show that collaboration modes are emergent, fine-grained, and dynamically reconstructed across time and context.
Together, our findings reframe collaboration success from static performance optimization to the ability to manage multi-objective trade-offs through dynamic adaptation. Parenting represents a high-stakes, relational, and asymmetric-agency context (Gurdal and Sorbring, 2018; Shin et al., 2021; Theofanopoulou and Slovak, 2022). In such contexts, parents must simultaneously assume responsibility while preserving children’s autonomy, which further intensifies the tensions among collaboration goals. Thus, our work also provides a new lens for understanding human-AI collaboration in other high-stakes, relational domains such as healthcare, education, and counseling (Saxena et al., 2021; Xhakaj et al., 2016; Borghoff et al., 2025).
7.4. Design Implications
Grounded in our findings, we derive design implications for flexible and effective parent–AI collaboration.
7.4.1. Context-aware Explainability without Disrupting Interaction
AI interventions in sensitive contexts can appear intrusive without explanation, undermining parental trust (Binns et al., 2018; Kulesza et al., 2015). Our results show that parents’ collaboration preferences shift systematically with emotional intensity, conversation stage, and parental goals. Systems should therefore provide context-aware explanations that clarify strategy–context links (Wang et al., 2019b; Dwivedi et al., 2023) in family communication settings.
However, continuous explanations risk cognitive overload and interrupt the smooth flow of parent-child conversation. Explainability should be treated not as a constant stream of information, but as an adaptive interactional feature, where timing, modality, and granularity are dynamically adjusted to balance transparency with conversational continuity.
7.4.2. Distinguishing Exploration from Stable Usage Patterns
Parents engaged with the system through an iterative process of trial and adjustment, often experimenting with new functions or revisiting previously dismissed ones. This means that system logs do not always reflect stable preferences. Therefore, when modeling user preferences (Pi et al., 2020; Jawaheer et al., 2014), systems should distinguish exploration from stable usage patterns.
Moreover, rather than treating exploration as noise, it should be recognized as a valuable meta-strategy that reflects how parents probe AI capabilities. Future systems may leverage such patterns to personalize support based on users’ exploration styles.
7.4.3. Rethinking AI Literacy in Parenting
Collaboration flexibility varied with parents’ AI literacy: in our study, high-literacy parents dynamically reconfigured functions, whereas low-literacy parents tended to rely on fixed collaboration modes. This suggests that parent–AI collaboration systems may impose excessive complexity and cognitive burden on low-literacy parents, highlighting the need for literacy-aware designs that adapt system complexity to parents’ capabilities.
However, in parenting contexts, AI literacy extends beyond technical proficiency (Long and Magerko, 2020; Ng et al., 2021) to include relational sensitivity—knowing when and how to involve AI without disrupting the parent–child relationship.
At the same time, overly simplified designs may constrain parents’ agency and hinder long-term learning. This underscores the importance of building on prior work on fostering AI literacy (Lee et al., 2021; Ng et al., 2024), and balancing immediate usability with the gradual development of parents’ AI literacy.
7.5. Limitations and Future Work
First, participants were limited to Chinese parents and children, with most parents being mothers, reflecting common caregiving patterns (McBride and Mills, 1993; Schoppe-Sullivan et al., 2013; Hernawati and Herawati, 2020). Future work should examine broader cultural contexts and parental roles, including mother–father differences in AI collaboration under similar structured interactions.
Second, the lab-based setting ensured controlled observation of moment-to-moment collaboration shifts but may not reflect everyday home communication routines. Longitudinal in-home deployments are needed to understand how parent–AI collaboration evolves in natural family settings.
Third, we did not assess children’s outcomes, as COMPASS was designed as a probe to surface parental strategies rather than a full intervention. Our contribution lies in theorizing and understanding parent-AI collaboration. Future research should further investigate how different parent–AI divisions of labor influence children’s communication and development.
Finally, COMPASS focused on six common scaffolding functions identified in our formative study, which do not cover the full design space. While this scope is sufficient for our purpose of examining parents’ shifting collaboration needs, future work could extend parent–AI collaboration to provide richer, more personalized support.
8. Conclusion
We developed COMPASS, a research probe designed to provide flexible scaffolding for parents during parent-child conversations. Using COMPASS as a probe in a study with 21 parent-child pairs, we find that parent–AI collaboration evolves from backstage support to partner-style co-performance, shifting with contextual factors to balance cognitive and emotional regulation. We identify three parental strategies, namely parent-oriented, child-oriented, and relationship-oriented strategies, extending Parental Goal Theory and informing the design of future parental support systems that enable flexible and effective partnerships.
Appendix A Details of the Six Parent Support Functions
This section presents the six AI support functions (F1-F6) identified from parents’ requests during the formative study.
Scenario Provision (F1). Parents preferred that the AI introduce topics through contextualized storytelling or illustrative examples to evoke children’s empathy and understanding, rather than directly providing fixed utterances. The ability of the AI to generate stories naturally aligns with everyday communication and children’s cognitive development stages, making it “easier for children to understand and accept” (P6).
Utterance Suggestion (F2). Seven out of nine parents expected the AI to provide concise, contextually appropriate utterance suggestions that they “could read aloud directly” (P8) at critical moments to help them overcome difficulties in phrasing. This highlights parents’ need for immediacy and contextual adaptability rather than reliance on generic templates.
Reflective Feedback (F3). Some parents emphasized their preference for heuristic rather than judgmental feedback: “I don’t want the AI to criticize me in a harsh tone… I need it to inspire me on how to think” (P2). Parents were more inclined to view the AI as a gentle thinking partner rather than an authoritative instructor, and therefore, the reflection function should focus on guidance and empathy.
Conversation Summarization (F4). Six parents reported that they often found it difficult to track their children’s emotions and perspectives, and thus hoped that the AI could summarize both parties’ positions and feelings in real time. One parent expressed a desire for the AI to analyze family members’ thoughts and reasoning, “so that everyone can understand each other, and I can respond to my child better” (P1). This reflects parents’ expectations for the AI to function as an information filter and emotional mediator, helping maintain smooth conversation and promote understanding.
Agent-as-Speaker (F5). Although most parents preferred the AI to act as an assistant, some welcomed its brief involvement as a “third-party friend” or “story character” when conversations stalled (P4). The premise was that “the AI should be a good listener”, intervening naturally and gently without interrupting parent–child interaction (P2).
Multimodal Enrichment (F6). Some parents noted that children aged 6-8 have difficulty understanding abstract vocabulary, and therefore suggested that the AI provide multimodal feedback such as picture book illustrations, enabling children to “be more willing to continue the conversation” (P5). This reflects parents’ desire for the system to reduce comprehension burden through concrete and vivid forms while maintaining engagement in the conversation.
Appendix B Implementation of the COMPASS Research Probe
We implemented the main system workflow on a Flask server, while the frontend interface is built in Vue and communicates with the server via WebSocket. The frontend runs on a touchscreen-enabled laptop to provide an intuitive and interactive interface for parents. For the Audio Processor, we utilized the TINGWU API, with audio sliced and compressed on the frontend prior to transcription. We also employed GPT-4o-mini to map each diarized speaker index to the corresponding role (parent or child). For all six functions in the Function Library, we leveraged the GPT-4o API to balance speed and output quality. For the Multimodal Enrichment Module, we employed the Stable Diffusion XL (SDXL) model due to speed considerations, whereas for the Agent-as-Speaker Module, we used Alibaba Cloud’s streaming synthesis API configured with a gentle female voice profile.
Appendix C Additional User Evaluation Details
This section provides participant demographic details for the user evaluation (Table 3).
Table 3. Participant demographic information in the user study.
ID
Child
Age
Child
Gender
Parent
Age
Parent
Gender
Parental
Education
1 6 Male 36 Female Bachelor’s
2 7 Male 38 Female Master’s
3 7 Male 43 Female Bachelor’s
4 8 Female 37 Female Bachelor’s
5 6 Female 42 Male Ph.D.
6 7 Female 32 Female Ph.D.
7 6 Female 33 Female Master’s
8 7 Male 45 Female Master’s
9 8 Female 38 Female Master’s
10 7 Male 41 Male Master’s
11 7 Female 38 Female Bachelor’s
12 8 Female 42 Female Bachelor’s
13 6 Male 47 Male Master’s
14 7 Male 33 Female Master’s
15 8 Male 42 Female Bachelor’s
16 7 Female 35 Female Master’s
17 6 Female 40 Female Master’s
18 7 Male 42 Male Master’s
19 6 Female 36 Female Bachelor’s
20 8 Male 37 Female Bachelor’s
21 8 Male 43 Female Associate’s